A 9 V battery is connected in series with a resistor .The terminal voltage is found to be 8 V.
Current through the circuit is measured as 5 A. What is the internal resistance of the battery?

Grade:12

1 Answers

Arun
5 years ago
As we know battery supplied voltage = E - Ir
Where E is the EMF of cell , I is the current passing through battery and r is internal resistance of battery . Putting all the values we get
8=9-5×r
=> r=1/5 =0.2 ohm